OpenStudy (anonymous):

A cylinder has a radius of 8 cm and a height of 14 cm. What is the volume of the cylinder? A. 112pi cm^3 b. 224pi cm^3 c. 896pi cm^3 d. 3584pi cm^3

OpenStudy (michele_laino):

here the volume V of a cylinder is giben by the subsequent formula: \[V = \pi \times {R^2} \times H\]

OpenStudy (mrnood):

ok - you don't need to work out the answer using pi - you can see that all the answers have pi left in so volume of a cylinder is th earea of the circular end times the height so area of circle = pi * r^2 so volume = pi * r^2 * h in your case r=8 h= 14 so put those number s into pi *r^2*h to get the right answer...

OpenStudy (michele_laino):

wher R is the radius, namely R= 8 cm, and H is the height of our cylinder, namely H=14 cm

OpenStudy (michele_laino):

so, substituting, we get: \[V = \pi \times {R^2} \times H = \pi \times {8^2} \times 14 = 896\pi \;{\text{c}}{{\text{m}}^{\text{3}}}\]

OpenStudy (anonymous):

@Michele_Laino ya this is the last question in the chapter practice, also did anyone else got booted off the site temp?

OpenStudy (er.mohd.amir):

volume of cylinder is v=pi*r*r*h where r is radius and is height